Matt's Eat 'N Run has two categories of direct labor: unskilled, which costs $20 per hour, and skilled, which costs $30 per hour. Management has established standards per "equivalent friendly meal," which has been defined as a typical meal consisting of a hamburger, a drink, and french fries. Standards have been set as follows. Skilled labor 3 minutes per equivalent meal Unskilled labor 7 minutes per equivalent meal For the year, Matt’s sold 190,000 equivalent friendly meals and incurred the following labor costs. Skilled labor 9,500 hours $ 291,000 Unskilled labor 11,000 hours 242,000 Required: a. Compute labor price and efficiency variances. b. Compute labor mix and yield variances.
Answer #1

Basic Calculations for standard costs

Basic Calculations
Standard hours for Actual Production (SH) = Standard time per unit x Actual Output
Standard hours for (Skilled Labor)
= 3 min x 190000 units / 60 min per hour
= 9500 Hours
Standard hours for (Unskilled Labor)
= 7 min x 190000 units / 60 min per hour
= 22166.67 Hours
Standard Mix Hour (or Revised Actual Hours) (SM) = Actual hours in standard Proportion
Standard Mix Hour (SM) (Skilled) = 20500 * 3 min / (3 min +7 Min)
Standard Mix Hour (SM) (Skilled) = 20500 * 3 min / (10 min)
Standard Mix Hour (SM) (Skilled) = 9500 Hours
Standard Mix Hour (SM) (Unskilled) = 20500 * 7 min / (3 min +7 Min)
Standard Mix Hour (SM) (Unskilled) = 20500 * 7 min / (10 min)
Standard Mix Hour (SM) (Skilled) = 22166.67 Hours

.

Key Data in standard format

Actual Production is 190000 units
Particulars Standard Actual Revised
Standard Hours Standard Rate per Hour Total Standard Cost Actual Hours Actual Rate per hour Total Actual Cost Standard Mix Hours  
SH SR SC AH AR AC SM
Labor
Skilled Labor 9500.00 30.00 285000.00 9500.00 30.63 291000.00 6150.00
Unskilled Labor 22166.67 20.00 443333.40 11000.00 22.00 242000.00 14350.00
Sub Total 31666.67 728333.40 20500.00 533000.00 20500.00

.

Calculation of Labor Variances.

(i) Labor Cost Variance = Standard Cost – Actual Cost
= (SR x SH) – (AR × AH)
= SC - AC
= 728333.4 - 533000
= $195,333.40 (Favorable)
(ii) Labor Price Variance = Standard Cost of Actual Time – Actual Cost
a) Labor Price Variance (Skilled Labor)
= (SR – AR) × AH
= (30.00 – 30.63) x 9500
= $5,985 (Unfavorable)
b) Labor Price Variance (Unskilled Labor)
= (SR – AR) × AH
= (20.00 – 22.00) x 11000
= $22,000 (Unfavorable)
c) Labor Price Variance (Total)
= (-5985) + 22000
= $16,015 (favorable)
(iii) Labor Efficiency Variance
= Standard Cost of Standard Time for Actual Production – Standard Cost of Actual Time
a) Labor Efficiency Variance (Skilled Labor)
= (SH – AH) × SR
= (9500 - 9500) x 30
= (0) x30
= Nil
b) Labor Efficiency Variance (Unskilled Labor)
= (SH – AH) × SR
= (22166.67 - 11000) x 20
= (11166.67) x20
= $223,333.40 (Favorable)
c) Labor Efficiency Variance (Total)
= 0 + 223333.40
= 223,333.40 (Favorable)
(iv) Labor Mix Variance = (Revised Standard Hours - Actual Hours) x Standard Rate
Labor Mix Variance = (SM - AH) x SR
a) Labor Mix Variance (Skilled)
= (6150 - 9500) x 30
= (-3350) x 30
= $100,500 (Unfavorable)
b) Labor Mix Variance (Unskilled)
= (14350-22166.67) x 20
= (-7816.67) x 20
= $156,333.40 (Unfavorable)
c) Labor Mix Variance (Total)
= (-100500) + (-156333.40)
= $256,833.40 (Unfavorable)
(v) Labor Yield Variance = ( Standard Hours - Revised Standard Hours) x Standard Rate
Labor Yield Variance = (SH - SM) x SR
a) Labor Yield Variance (Skilled)
= (9500 - 6150) x 30
= (3350) x 30
= $100,500 (Favorable)
b) Labor Yield Variance (Unskilled)
= (22166.67 - 14350) x 20
= (7816.67) x 20
= $156,333.40 (Favorable)
c) Labor Yield Variance (Total)
= 100,500 + 156,333.40
= 256,833.40 (Favorable)